y''-3y'+2y=2x2+ex+2xex+4e3x

1
Expert's answer
2021-08-09T16:52:14-0400

"\displaystyle
\textsf{The auxiliary equation is}\\
m^2 - 3m + 2 = 0\\
m = 1, 2\\

\implies y = C_1e^x + C_2e^{2x}\\

y''-3y'+2y=2x^2+e^x+2xe^x+4e^(3x)\\

y = Ax^2 + Bx + C + Dxe^x + Ex^2e^x + Fe^{3x}\\
y' = 2Ax + B + De^x + Dxe^x + 2Exe^x + Ex^2e^x + 3Fe^{3x}\\
\begin{aligned}
y'' &= 2A + 2De^x + Dxe^x + 2Ee^x + 2Exe^x \\&+ Ex^2e^x + 2Exe^x + 9Fe^{3x}
\\&= 2A + (2D + 2E)e^x + (D + 4E)xe^x + Ex^2e^x + 9Fe^{3x}
\end{aligned}\\

y' = 2Ax + B + De^x + (D + 2E)xe^x + Ex^2e^x + 3Fe^{3x}\\
y = Ax^2 + Bx + C + Dxe^x + Ex^2e^x + Fe^{3x}\\

\begin{aligned}
y" - 3y' + 2y &= (2A - 3B + 2C) + (2E - D)e^x - 2Exe^x \n\\\\&+ 2Fe^{3x} + (2B - 6A)x + 2Ax^2 \n\\\\&= 2x^2+e^x+2xe^x+4e^(3x)\n\\end{aligned}\\\\\n\n2A = 2, A = 1\\\\\n\n2B - 6A = 0, B = 3\\\\\n\n2F = 4, F = 2\\\\\n\n-2E = 2, E = -1\\\\\n\n2E - D = 1, D = -3\\\\\n\n2A - 3B + 2C = 0\\\\\n\n2 - 9 + 2C = 0, C = 7\/2\\\\\n\ny = x^2 + 3x + 7\/2 - 3xe^x - x^2e^x + 2e^{3x}\\\\\n\n\\implies y = C_1e^x + C_2e^{2x} + x^2 + 3x \\\\+ 7\/2 - 3xe^x - x^2e^x + 2e^{3x}"